OpenStudy (anonymous):

Find the square root that is a real number. √-9

OpenStudy (anonymous):

no real number 3i

OpenStudy (anonymous):

3i

OpenStudy (anonymous):

3i is not real number

OpenStudy (anonymous):

3i

OpenStudy (anonymous):

i is imagery number...

OpenStudy (anonymous):

0+3i

OpenStudy (anonymous):

you're question is incorrect. 3i is the answer, but it is not real.

OpenStudy (anonymous):

\[\sqrt{9}\times \sqrt{-1} = 3\sqrt{-1}\] \sqrt{-1}\] = Imagery number... So, there is no real number.

OpenStudy (anonymous):

THX

OpenStudy (anonymous):

Yw.. ^.^

OpenStudy (anonymous):

Question is correct, and answer is it's impossible to find :)